OpenStudy (anonymous):

CHECK WORK? X 11 14 16 19 21 23 24 29 P(X) 0.07 0.21 0.17 0.25 0.05 0.04 0.1… The mean of the discrete random variable X is 18.59. What is the variance of X? Round your answer to the nearest hundredth. (I got 31.132) Var(X) = (11² x 0.07) + (14² x 0.21) + (16² x 0.17) + (19² x 0.25) + (21² x 0.05) + (23² x 0.04) + (24² x 0.1) + (29² x 0.11) Var(X) = 376.72 - 18.59²

OpenStudy (anonymous):

P(X)0.07,0.21,0.17,0.25,0.05,0.04,0.13,0.08

OpenStudy (anonymous):

mmm I don't believe that's quite right.

OpenStudy (anonymous):

Yea i just seen that..

OpenStudy (anonymous):

For each X do: P(X)(difference from the mean)^2

OpenStudy (anonymous):

That is, P(X)(X-18.59)^2

OpenStudy (anonymous):

an then add them up an divide by 8?

OpenStudy (anonymous):

Add them up. Don't divide in this case.

OpenStudy (anonymous):

When it's equally likely chances, we divide by the number of possibilities. When each possibility has a different probability of happening, we just multiply each difference by that probability, but we don't have to divide by the number of possibiliities.

OpenStudy (anonymous):

Alright,. well I did it, got 267.8929 rounded.... 267.893

OpenStudy (anonymous):

Ahhh something is wrong. Here is what you should have calculated for the first X, X=11: P(X)*(X-18.59)^2 = 0.07*(11-18.59)^2 = 4.032567

OpenStudy (anonymous):

ahh i missed multiplying the P(X)

OpenStudy (anonymous):

Important step =)

OpenStudy (anonymous):

I hate math... its to time consuming.

OpenStudy (anonymous):

It's usually not so bad. But yeah, this kind of problem is tedious. May I suggest using www.wolframalpha.com?

OpenStudy (anonymous):

It's pretty quick and easy to use, can do a lot more than your calculator, and if you make a mistake, you can just go back and edit what you messed up on instead of starting over completely.

OpenStudy (anonymous):

Plus, you can copy and paste the link of what you get, and I can look over it and figure out where you went wrong.